OpenStudy (anonymous):

Find the distance between the given points. A(5, 3) and B(7, -4)

OpenStudy (anonymous):

= sqrt[ (7-5)^2 + (3--4)^2] = sqrt 53 = 7.28

OpenStudy (anonymous):

think with a triangle \[D=\sqrt{(7-5)²+(-4-3)²}\]\[ D=\sqrt{4+49}\]\[ D=\sqrt{53}\]

OpenStudy (anonymous):

the distance formula is: \[\sqrt{(y_{2}-y_{1})^{2}+(x_{2}-x_{1})^{2}}\]so use point A as 1 and point B as 2 you will get: \[\sqrt{(7-5)^{2}+(-4-3)^{2}}=\sqrt{4+49}=\sqrt{53}\]